# Adaptive Self-Retention & Correction for Continual Learning

### v2

@@ Abstract
Continual learning, also known as lifelong learning or incremental learning, refers to the process by which a model learns from a stream of incoming data over time. A common problem in continual learning is the classification layer’s bias towards the most recent task. Traditionally, methods have relied on incorporating data from past tasks during training to mitigate this issue. However, the recent shift in continual learning to memory-free environments has rendered these approaches infeasible. In this study, we propose a solution focused on the testing phase. We first introduce a simple Out-of-Task Detection method, OTD, designed to accurately identify samples from past tasks during testing. Leveraging OTD, we then propose: (1) an Adaptive Self-Retention mechanism for dynamically tuning the classifier layer on past task data; (2) an Adaptive Self-Correction mechanism for revising predictions when the model classifies data from previous tasks into classes from the current task. We name our approach Adaptive Self-Retention & Correction (ARC). While designed for memory-free environments, ARC also proves effective in memory-based settings. Extensive experiments show that our proposed method can be plugged in to virtually any existing continual learning approach without requiring any modifications to its training procedure. Specifically, when integrated with state-of-the-art approaches, ARC achieves an average performance increase of 2.7% and 2.6% on the CIFAR-100 and Imagenet-R datasets, respectively.
@@ 1 Introduction
Deep learning has seen remarkable advancements over the past decade (Krizhevsky et al., 2012, Ren et al., 2015, He et al., 2016, Ronneberger et al., 2015, Long et al., 2015, Dosovitskiy et al., 2021), largely due to the availability of large static datasets. However, real-life applications often require continually updating models on new data. A major challenge in this dynamic updating process is the phenomenon of catastrophic forgetting (Kirkpatrick et al., 2016, French, 1999), where a model’s performance on previously learned tasks drastically degrades. To address this issue, various continual learning (CL) methods (De Lange et al., 2021, Hadsell et al., 2020, Mai et al., 2022) have been developed to provide a more sustainable and efficient approach to sequential model adaptation.
The community has identified two primary sources of catastrophic forgetting: 1) overfitting of the feature extraction network on new tasks (Li and Hoiem, 2017, Kirkpatrick et al., 2016, Rebuffi et al., 2017); and 2) bias towards new classes introduced by the linear classifier (Wu et al., 2019, Chrysakis and Moens, 2023, Belouadah and Popescu, 2019). Throughout the years, most works have been focusing on addressing forgetting of the representation layer. More recently, however, Zhang et al. (2023) show that for pretrained models, using a small learning rate for the representation layer and a larger learning rate for the classifier can effectively mitigate overfitting issue, thereby relieving catastrophic forgetting. In this work, we take a step further and show that even for models overfitted on the latest task, the amount of information lost on the representation layer is minimal. As such, we primarily focus on tackling the latter issue, i.e., mitigating classifier bias.
@@
With these two categories of predictions, we design two independent approaches tailored to each. Specifically, we propose Adaptive Self-Retention & Correction (ARC) which consists of:
1. Adaptive Self-Retention, where we utilize accurately classified samples from previous tasks and re-balance the classification layer in an online manner with one gradient update. 2. Adaptive Self-Correction, where we correct predictions made on potentially misclassified samples from previous tasks by adjusting the classifier’s preference for the most recent task’s classes.
@@
    We leverage the fact that models can interact with samples from past tasks during inference and propose Adaptive Self-Retention & Correction (ARC), where the classifier weight is re-balanced and predictions on potentially misclassified images are rectified.
@@ 3.3 Out-of-Task Detection
3.4 Adaptive Self-Retention & Correction
Based on OTD, we design two independent approaches, Adaptive Self-Retention and Adaptive Self-Correction, that leverage the distinct characteristics of the data following 1 and 2 to mitigate classifier bias. Together, they form the proposed Adaptive Self-Retention & Correction approach. An overview of ARC is shown in Algorithm 1.
Algorithm 1 Adaptive Self-Retention & Correction (ARC)
Inputs: Current task identity $t$; Incremental step size $s$; Testing image $\bm{x}$; Initially predicted class $\hat{y}$; Adaptive Self-Retention threshold $\beta$; Adaptive Self-Correction threshold $\gamma$; Task-based Softmax Score temperature $T$;network $M_{\theta}(\cdot)=h_{\theta_{cls}}(f_{\theta_{rps}}(\cdot))$ with parameters $\theta=\{\theta_{rps},\theta_{cls}\}$;
@@ 3.4 Adaptive Self-Retention & Correction
3.4.1 Adaptive Self-Retention
We first introduce Adaptive Self-Retention using data following 1. Traditional methods for reducing classifier bias mainly rely on training with data from the past task and are not applicable in memory-free situations. However, by using OTD, we can still leverage data from past tasks, but this time adaptively during the testing phase. To do so, we utilize the predicted labels as supervisory signals and train the classifier with the backbone network frozen using a cross-entropy loss. In this way, the tendency within the classifier towards classes of the current task can be effectively mitigated by re-balancing weights towards previous classes. Furthermore, since predicted labels inevitably introduce errors and uncertainty, we employ an additional entropy minimization loss. This approach enhances the model’s confidence in its predictions, thereby offsetting the adverse effects of noise and ensuring more reliable outputs during testing. Finally, to align more closely with practical applications where test samples arrive in an online manner, we choose to train the classifier with only one gradient update on each sample.
@@ 4.2 Combining ARC with State-of-the-Art Continual Learning Methods
Comparison with TTA method. Since Test Time Adaptation (TTA) is a special type of model correction, we also report performance comparison with a representative TTA method, TENT (Wang et al., 2020) in Table 3. As demonstrated, while TENT exhibits some performance improvements, the magnitude is consistently lower compared to ARC. Specifically, on Split CIFAR100 Inc10, TENT’s average performance gains are 1.8% lower than those of ARC, and on Split ImageNet-R Inc20, the gap increases to 2.9%. We hypothesize that there are two primary reasons for this. First, TENT, and similar TTA methods, adapt to all incoming test data, whereas our approach selectively adapts only to past task data. Second, TENT’s adaptation is comparable to Adaptive Self-Retention but lacks the ability to identify and correct potentially misclassified data from previous tasks, i.e., Adaptive Self-Correction, further limiting its effectiveness.
@@ 4.3 In-Depth Analysis of ARC and Its Properties
Detailed contribution. We further present the specific contributions of Adaptive Self-Retention and Adaptive Self-Correction  individually in Table 4. The average increase using Adaptive Self-Retention is 1.8% on Split CIFAR-100 and Split Imagenet-R, and the average increase using Adaptive Self-Correction is 1.1% on Split CIFAR-100 and 1.2% on Split Imagenet-R. Notably, the performance boost when applying Adaptive Self-Correction to prompt-based methods is low. We find that the small size of this boost is due to the small number of test samples for these prompt based methods that satisfy 2. For example, on Split Imagenet-R, only 1.9% of test samples satisfy 2 for L2P, while for iCarL, 7.4% of test samples satisfy this assumption.
@@
Table 4: Specific contribution of Adaptive Self-Retention and Adaptive Self-Correction.
@@
((a)) Ablation for Adaptive Self-Retention.
@@
((b)) Ablation for Adaptive Self-Correction.
Figure 4: Ablation studies for specific components of Adaptive Self-Retention and Adaptive Self-Correction on Split CIFAR-100.
@@ 4.4 Ablation Studies
Adaptive Self-Retention. We present in Figure 4(a) the ablation studies on the objective of the training process of Adaptive Self-Retention, where the effect of the cross-entropy loss $\mathcal{L}_{CE}$ and the entropy minimization loss $\mathcal{L}_{EM}$ on the Split CIFAR-100 dataset is tested. As is shown, while for L2P, the use of $\mathcal{L}_{CE}$ results in a slightly worse performance, in general, both are important components of Adaptive Self-Retention.
Adaptive Self-Correction. We further present in Figure 4(b) the ablation studies on our design of Adaptive Self-Correction. We first validate the necessity of the scaling temperature value for TSS. We see that when it is excluded, Finetune, iCarL, Der and Memo would experience an average performance drop of 0.6%. We also test our design of $w$ for 2. Specifically, we examine whether our concern over the two possible outcomes of low-confidence prediction is necessary. To do so, we compared against letting $w=c\leq\gamma$, i.e., making $w$ represent the raw confidence for the predicted class. As is demonstrated, such change would result in an average performance degradation of 0.5%, indicating the success of our specific design of $w$.
@@ 5 Conclusion
In this study, we addressed the critical issue of classifier bias in continual learning, especially in the current paradigm of leveraging pretrained models under memory-free settings, during the testing phase. We first built an Out-of-Task detection method, where we identified two special cases of testing data: 1) data from previous tasks that is correctly classified and 2) data from previous tasks that is misclassified to the current task. Consequently, we proposed two independent approaches, Adaptive Self-Retention and Adaptive Self-Correction, specifically tailored for each case. Together they form our approach Adaptive Self-Retention & Correction (ARC). Adaptive Self-Retention focuses on dynamically tuning the classifier layer and Adaptive Self-Correction implements a training-free bias correction mechanism. By leveraging the ability to access past task data during inference, ARC effectively mitigates classifier bias without requiring any modifications to the training procedure. Extensive experiments show that integrating ARC to existing methods can significantly boost their overall performance, demonstrating the effectiveness and robustness of our method.
@@
In this section, we provide further details to ensure the reproducibility of our work. Specifically, for Adaptive Self-Retention, given a batch of test samples, we compute the loss using Equation 3 and perform a single gradient update based on this loss. For non-prompt-based methods, we utilize the SGD optimizer, while for prompt-based methods, we employ the Adam optimizer.

### v4

# Adaptive Retention & Correction: Test-Time Training for Continual Learning

@@ Abstract
Continual learning, also known as lifelong learning or incremental learning, refers to the process by which a model learns from a stream of incoming data over time. A common problem in continual learning is the classification layer’s bias towards the most recent task. Traditionally, methods have relied on incorporating data from past tasks during training to mitigate this problem. However, the recent shift in continual learning to memory-free environments has rendered these approaches infeasible. In this study, we propose a solution focused on the testing phase. We first introduce a simple Out-of-Task Detection method, OTD, designed to accurately identify samples from past tasks during testing. Using OTD, we then propose: (1) an Adaptive Retention mechanism for dynamically tuning the classifier layer on past task data; (2) an Adaptive Correction mechanism for revising predictions when the model classifies the data from previous tasks into classes of the current task. We name our approach Adaptive Retention & Correction (ARC). Although designed for memory-free environments, ARC also proves effective in memory-based settings. Extensive experiments show that our proposed method can be plugged in to virtually any existing continual learning approach without requiring any modifications to its training procedure. Specifically, when integrated with state-of-the-art approaches, ARC achieves an average performance increase of 2.7% and 2.6% on the CIFAR-100 and Imagenet-R datasets, respectively. Code is available at Github Link. Corresponding author.
@@ 1 Introduction
Deep learning has seen remarkable advances over the past decade (Krizhevsky et al., 2012, Ren et al., 2015, He et al., 2016, Ronneberger et al., 2015, Long et al., 2015, Dosovitskiy et al., 2021), largely due to the availability of large static datasets. However, real-life applications often require continually updating models on new data. A major challenge in this dynamic updating process is the phenomenon of catastrophic forgetting (Kirkpatrick et al., 2016, French, 1999), where a model’s performance on previously learned tasks drastically degrades. To address this issue, various continual learning (CL) methods (De Lange et al., 2021, Hadsell et al., 2020, Mai et al., 2022) have been developed to provide a more sustainable and efficient approach to sequential model adaptation.
The community has identified two primary sources of catastrophic forgetting: 1) overfitting of the feature extraction network on new tasks (Li and Hoiem, 2017, Kirkpatrick et al., 2016, Rebuffi et al., 2017); and 2) bias towards new classes introduced by the linear classifier (Wu et al., 2019, Chrysakis and Moens, 2023, Belouadah and Popescu, 2019). Throughout the years, most works have been focusing on addressing forgetting of the representation layer. More recently, however, Zhang et al. (2023) show that for pretrained models, using a small learning rate for the representation layer and a larger learning rate for the classifier can effectively mitigate the overfitting issue, thereby relieving catastrophic forgetting. In this work, we take a step further and show that even for models overfitted on the latest task, the amount of information lost on the representation layer is minimal. As such, we primarily focus on tackling the latter issue, i.e., mitigating classifier bias.
@@
With these two categories of predictions, we design two independent approaches tailored to each. Specifically, we propose Adaptive Retention & Correction (ARC) which consists of:
1. Adaptive Retention, where we utilize accurately classified samples from previous tasks and re-balance the classification layer in an online manner with one gradient update. 2. Adaptive Correction, where we correct predictions made on potentially misclassified samples from previous tasks by adjusting the classifier’s preference for the most recent task’s classes.
@@
    We leverage the fact that models can interact with samples from past tasks during inference and propose Adaptive Retention & Correction (ARC), where the classifier weight is re-balanced and predictions on potentially misclassified images are rectified.
@@ 3.3 Out-of-Task Detection
3.4 Adaptive Retention & Correction
Based on OTD, we design two independent approaches, Adaptive Retention and Adaptive Correction, that leverage the distinct characteristics of the data following 1 and 2 to mitigate classifier bias. Together, they form the proposed Adaptive Retention & Correction approach. An overview of ARC is shown in Algorithm 1.
Algorithm 1 Adaptive Retention & Correction (ARC)
Inputs: Current task identity $t$; Incremental step size $s$; Testing image $\bm{x}$; Initially predicted class $\hat{y}$; Adaptive Retention threshold $\beta$; Adaptive Correction threshold $\gamma$; Task-based Softmax Score temperature $T$;network $M_{\theta}(\cdot)=h_{\theta_{cls}}(f_{\theta_{rps}}(\cdot))$ with parameters $\theta=\{\theta_{rps},\theta_{cls}\}$;
@@ 3.4 Adaptive Retention & Correction
3.4.1 Adaptive Retention
We first introduce Adaptive Retention using data following 1. Traditional methods for reducing classifier bias mainly rely on training with data from the past task and are not applicable in memory-free situations. However, by using OTD, we can still leverage data from past tasks, but this time adaptively during the testing phase. To do so, we utilize the predicted labels as supervisory signals and train the classifier with the backbone network frozen using a cross-entropy loss. In this way, the tendency within the classifier towards classes of the current task can be effectively mitigated by re-balancing weights towards previous classes. Furthermore, since predicted labels inevitably introduce errors and uncertainty, we employ an additional entropy minimization loss. This approach enhances the model’s confidence in its predictions, thereby offsetting the adverse effects of noise and ensuring more reliable outputs during testing. Finally, to align more closely with practical applications where test samples arrive in an online manner, we choose to train the classifier with only one gradient update on each sample.
@@ 4.2 Combining ARC with State-of-the-Art Continual Learning Methods
Comparison with TTA method. Since Test Time Adaptation (TTA) is a special type of model correction, we also report performance comparison with a representative TTA method, TENT (Wang et al., 2020) in Table 3. As demonstrated, while TENT exhibits some performance improvements, the magnitude is consistently lower compared to ARC. Specifically, on Split CIFAR100 Inc10, TENT’s average performance gains are 1.8% lower than those of ARC, and on Split ImageNet-R Inc20, the gap increases to 2.9%. We hypothesize that there are two primary reasons for this. First, TENT, and similar TTA methods, adapt to all incoming test data, whereas our approach selectively adapts only to past task data. Second, TENT’s adaptation is comparable to Adaptive Retention but lacks the ability to identify and correct potentially misclassified data from previous tasks, i.e., Adaptive Correction, further limiting its effectiveness.
@@ 4.3 In-Depth Analysis of ARC and Its Properties
Detailed contribution. We further present the specific contributions of Adaptive Retention and Adaptive Correction  individually in Table 4. The average increase using Adaptive Retention is 1.8% on Split CIFAR-100 and Split Imagenet-R, and the average increase using Adaptive Correction is 1.1% on Split CIFAR-100 and 1.2% on Split Imagenet-R. Notably, the performance boost when applying Adaptive Correction to prompt-based methods is low. We find that the small size of this boost is due to the small number of test samples for these prompt based methods that satisfy 2. For example, on Split Imagenet-R, only 1.9% of test samples satisfy 2 for L2P, while for iCarL, 7.4% of test samples satisfy this assumption.
@@
Table 4: Specific contribution of Adaptive Retention and Adaptive Correction.
@@
((a)) Ablation for Adaptive Retention.
@@
((b)) Ablation for Adaptive Correction.
Figure 4: Ablation studies for specific components of Adaptive Retention and Adaptive Correction on Split CIFAR-100.
@@ 4.4 Ablation Studies
Adaptive Retention. We present in Figure 4(a) the ablation studies on the objective of the training process of Adaptive Retention, where the effect of the cross-entropy loss $\mathcal{L}_{CE}$ and the entropy minimization loss $\mathcal{L}_{EM}$ on the Split CIFAR-100 dataset is tested. As is shown, while for L2P, the use of $\mathcal{L}_{CE}$ results in a slightly worse performance, in general, both are important components of Adaptive Retention.
Adaptive Correction. We further present in Figure 4(b) the ablation studies on our design of Adaptive Correction. We first validate the necessity of the scaling temperature value for TSS. We see that when it is excluded, Finetune, iCarL, Der and Memo would experience an average performance drop of 0.6%. We also test our design of $w$ for 2. Specifically, we examine whether our concern over the two possible outcomes of low-confidence prediction is necessary. To do so, we compared against letting $w=c\leq\gamma$, i.e., making $w$ represent the raw confidence for the predicted class. As is demonstrated, such change would result in an average performance degradation of 0.5%, indicating the success of our specific design of $w$.
@@ 5 Conclusion
In this study, we addressed the critical issue of classifier bias in continual learning, especially in the current paradigm of leveraging pretrained models under memory-free settings, during the testing phase. We first built an Out-of-Task detection method, where we identified two special cases of testing data: 1) data from previous tasks that is correctly classified and 2) data from previous tasks that is misclassified to the current task. Consequently, we proposed two independent approaches, Adaptive Retention and Adaptive Correction, specifically tailored for each case. Together they form our approach Adaptive Retention & Correction (ARC). Adaptive Retention focuses on dynamically tuning the classifier layer and Adaptive Correction implements a training-free bias correction mechanism. By leveraging the ability to access past task data during inference, ARC effectively mitigates classifier bias without requiring any modifications to the training procedure. Extensive experiments show that integrating ARC to existing methods can significantly boost their overall performance, demonstrating the effectiveness and robustness of our method.
@@
In this section, we provide further details to ensure the reproducibility of our work. Specifically, for Adaptive Retention, given a batch of test samples, we compute the loss using Equation 3 and perform a single gradient update based on this loss. For non-prompt-based methods, we utilize the SGD optimizer, while for prompt-based methods, we employ the Adam optimizer.